OpenStudy (anonymous):

i need a little help . . Solve for elimination -4x-9y=-2 -9x-18y=0

sam (.sam.):

-4x-9y=-2 times 2 -9x-18y=0 -------------------- -8x-18y=-4 -9x-18y=0 x =-4 -------------------- -9x-18y=0 y=-2

Parth (parthkohli):

1) Multiply the top equation(both sides) by 2 2) Add both equations. 3) Solve for x. 4) Substitute the value of x to find the the value of y.

OpenStudy (anonymous):

thanks so much . saved my life ^__^

sam (.sam.):

yw

OpenStudy (callisto):

Wait... y is 2 ... not -2

OpenStudy (callisto):

-9x-18y=0 -9(-4)-18y=0 -18y = -36 y=2 ....

OpenStudy (anonymous):

ok , thanks for finalizing that .